# MimicMotion: High-Quality Human Motion Video Generation with Confidence-aware Pose Guidance

### v1

@@ 3.4 Confidence-aware pose guidance
(a) The occluded areas have lower confidence.
(b) The motion blurred areas have lower confidence.
Figure 3: Pose estimation shows uncertainty in some areas with occlusion or dynamic blurring caused by body motions. The confidence score is lower in these areas. We convey this information to Unet.
Inaccurate pose estimation has a negative impact on the model’s training and inference. Accurately estimating poses from images is challenging in dynamic videos. Estimating poses from 2D images is inherently difficult. The limited capability of the pose estimation model, like DWPose [35], is only one aspect of the reason. The more significant cause is the inherent uncertainty of pose from dynamic appearances and motions. Specifically, incorrect pose guidance signals can mislead the model, resulting in the generation of inaccurate or distorted outputs, as illustrated in Figure 9. Moreover, noisy pose guidance signals can lead to overfitting on samples with incorrect poses, potentially causing training instability. This in turn may hinder the model’s ability to benefit from extended training schedules.
For this problem, we propose confidence-aware pose guidance, which leverages the confidence scores associated with each keypoint from the pose estimation model. These scores reflect the likelihood of accurate detection, with higher values indicating higher visibility, less occlusion, and motion blur. Instead of applying a fixed confidence threshold to filter the keypoints, as commonly adopted in prior works [38, 4], we utilize brightness on the pose guidance frame to represent the confidence level of pose estimation. Specifically, we integrate the confidence scores of the pose and keypoints into their respective drawing colors. This means that we multiply the color assigned to each keypoint and limb by its confidence score. Consequently, keypoints and corresponding limbs with higher confidence scores will appear more significant on the pose guidance map. This method enables the model to prioritize more reliable pose information in its guidance, thereby enhancing the overall accuracy of pose-guided generation.
Figure 3 illustrates this concept, showing how confidence-aware pose frames reflect situations of occlusion and motion blur. In this way, the uncertainty of pose estimation can be conveyed through the pose guidance, making pose guidance more informative. Our ablation studies show the effectiveness of this technique in suppressing visual artifacts, as shown in Figure 9.
@@
Moreover, we employ pose estimation and the associated confidence scores to alleviate region-specific artifacts, such as hand distortion, which are prevalent in the diffusion-based image and video generation models. Specifically, we identify reliable regions via thresholding keypoint confidence scores. By setting a threshold, we can distinguish between keypoints that are confidently detected and those that may be ambiguous or incorrect due to factors like occlusion or motion blur. Keypoints with confidence scores above the threshold are considered reliable. We implement a masking strategy that generates masks based on a confidence threshold. We unmask areas where confidence scores surpass a predefined threshold, thereby identifying reliable regions. When computing the loss of the video diffusion model, the loss values corresponding to the unmasked regions are amplified by a certain scale so they can have more effect on the model training than other masked regions.
@@
Specifically, to mitigate hand distortion, we compute masks using a confidence threshold for keypoints in the hand region. Only hands with all keypoint confidence scores exceeding this threshold are considered reliable, as a higher score correlates to higher visual quality. We then construct a bounding box around the hand by padding the boundary of these keypoints, and the enclosed rectangle is designated as unmasked. This region is subsequently assigned a larger weight in the loss calculation during the training of the video diffusion model. This selective unmasking and weighting process biases the model’s learning towards hands, especially hands with higher visual quality, effectively reducing distortion and improving the overall realism of the generated content.
3.5 Progressive latent fusion for long video generation
@@
Figure 4: This is an overview of our approach for long video generation. The latent features of video frames are represented with colored boxes. The darkness of color means the weight in latent fusion and darker color means a heavier weight. The dashed boxes represent video frame features involved in latent fusion.
@@
Limited by computation resources, generating long videos containing a large number of frames is challenging. For this problem, latent fusion during denoising with DM has been validated by some prior works like MultiDiffusion [33] which utilizes latent fusion on overlapped tiles to realize panoramic image generation. A similar idea can be applied to the video generation task. A straightforward approach is directly applying MultiDiffusion in the time domain, as in Lumiere [34]. Compared with spatial discontinuity between image tiles, viewers are more sensitive to temporal discontinuity because they can significantly cause flickering or even abrupt changes in content. For this problem, we propose a progressive approach for generating long videos with high temporal continuity.
@@
Progressive latent fusion is training-free and is integrated into the denoising process of the latent diffusion model during inference. Figure 4 shows an overview of this process. We omit the VAE for brevity. The denoising process is done in latent space in our method. In general, there are $T$ denoising steps in total and our latent fusion is applied within each step. For a long given pose sequence, we use a pre-defined strategy for splitting the whole sequence into segments, consisting of a fixed number of frames per segment (denoted as $N$), with a certain number ($C$) of overlapped frames between every two adjacent segments. For the sake of generation efficiency, it is common to assume that $C\ll N$. During each denoising step, video segments are firstly denoised separately with the trained model, conditioning on the same reference image and the corresponding sub-sequence of poses. Algorithm 1 shows the specific details of progressive latent fusion. As inputs, the reference image is denoted as $I_{\text{ref}}$ (c.f. Sec. 3.2) and pose frames corresponding to $j$-th frame in $i$-th video segment is denoted as $P_{i}^{j}$. We use $\mathbf{z}_{i}^{j}$ to denote the latent feature of $j$-th frame in $i$-th video segment. The denoising process starts from the maximum time step $T$ and the latent features are initialized with a normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Within each denoising step at time step $t$, the reversed diffusion process defined by the trained model (DM) is applied to the latent features of each video segment numbered $i$ separately, with $\mathbf{z}_{i}$, $I_{\text{ref}}$, $P_{i}$ and $t$ as inputs. During the latent fusion stage, for every two adjacent video segments, the involved video frames are then fused. To avoid the corruption of temporal smoothness near video segment boundaries after latent fusion, we propose progressive latent fusion. For a video frame involved in latent fusion, its fusion weight is determined by its relative position in the video segment it belongs to. Specifically, if a frame is close to the segment it belongs to, it will be assigned a heavier weight. For implementation, a fusion scale is pre-defined as $\lambda_{\text{fusion}}=1/(C+1)$ for controlling the level of latent fusion.
Input: $I_{\text{ref}}$: Reference image. $P_{i}^{j}$: Pose frame corresponding to $j$-th frame in $i$-th video segment; $\mathbf{z}_{i}^{j}$: The latent feature of $j$-th frame in $i$-th video segment; $N$: the number of frames in a video segment; $C$: the number of overlapped frames.
Output: $\mathbf{z}^{\prime}$: A long sequence of latent features of video frames.
$\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ ;
  // Random initialization of video latent features.
$\lambda_{\text{fusion}}\leftarrow 1/(C+1)$ ;
  // Set a scale of latent fusion.
for $t=T$ to $1$ do // Denoising from noisy latent features step by step.
       for $i=1,2,\dots$ do
            $\mathbf{z}_{i}\leftarrow\text{DM}(\mathbf{z}_{i},I_{\text{ref}},P_{i},t)$ // Separately denoise each segment.
      for $i=1,2,\dots$ do // Within each video segment.
             for $j=1$ to $N$ do // Start latent fusion for each frame.
                   if $i>1$ and $j\leq C$ then // Latent fusion with the previous segment.
                         $\mathbf{z}_{i}^{j}\leftarrow j\lambda_{\text{fusion}}\mathbf{z}_{i}^{j}+(1-j\lambda_{\text{fusion}})\mathbf{z}_{i-1}^{N-C+j}$
                   else if $j>N-C$ then // Latent fusion with the next segment.
                         $\mathbf{z}_{i}^{j}\leftarrow(N+1-j)\lambda_{\text{fusion}}\mathbf{z}_{i}^{j}+(C-N+j)\lambda_{\text{fusion}}\mathbf{z}_{i+1}^{C-N+j}$
return $\mathbf{z}^{\prime}=\text{Merge}(\mathbf{z})$ ;
  // Merge multi-segment features following Listing LABEL:lst:merge_function.
Algorithm 1 Progressive frame-level latent fusion for long video generation.
After applying $T$ iteratively denoising steps, a merging strategy denoted as Merge is used to get the final long sequence based on the denoised overlapped video segments in latent space. The function of Merge concatenates the multi-segment latents, which is described in Listing LABEL:lst:merge_function in detail.
Listing 1: Merge is for merging 2D list z representing overlapped video segments into a long list zp.
def Merge(z, C):
zp = z[0]
for i in range(1, len(z)):
zp.extend(z[i][C:])
return zp
@@ 4.1 Implementation details
We train our model on 4,436 human dancing videos collected from the internet. The average length of the training videos is 20.1s. We adopt the pre-trained weights from the public stable video diffusion 1.1 image-to-video model. The PoseNet is trained from scratch. We train our model on 8 NVIDIA A100 GPUs (40G) for 20 epochs, with a per-device batch size of 1. The learning rate is $10^{-5}$ with a linear warmup for the first 500 iterations. We tune all the parameters in the UNet and PoseNet.
@@ 4.2 Comparison to state-of-the-art methods
We compare our method with latest state-of-the-art pose-guided human video generation methods, including MagicPose [4], Moore-AnymateAnyone [38], and MuseV [39]. Following previous works [3, 4], we adopt the TikTok [40] dataset and use sequence 335 to 340 for testing.
We provide both qualitative and quantitative comparisons, complemented by a user study. Each method has a different input aspect ratio. To ensure a fair comparison, we only consider the central square region of the videos. Specifically, to accommodate each method’s unique input aspect ratios, we individually apply a center crop to the reference image and pose sequence. Then, we extract the center squares from the generated videos for a fair comparison across different methods. This applies to all experiments in comparison to state-of-the-art methods.
@@
Figure 5: Qualitative comparison to the state-of-the-art methods on TikTok test set. Our method achieves better hand generation quality, and more accurately adheres to the reference pose. Note that our method is not trained on the TikTok dataset.
Figure 6: Comparison of temporal smoothness with state-of-the-art methods. We visualize the 106th frame from seq 338 of the TikTok dataset and the pixel-wise difference between consecutive frames. MagicPose exhibits abrupt transitions, while Moore and MuseV show instability in texture and text. In contrast, our method demonstrates stable inter-frame differences and better temporal smoothness.
Importantly, our method shows superior temporal smoothness, characterized by smooth motion and minimal flickering. To illustrate this aspect, we present the pixel-wise differences between consecutive frames in Figure 6, which effectively illustrate the temporal stability of our method. From the figure, it is evident that MagicPose [4] exhibits abrupt transitions, Moore-AnymateAnyone [38] shows flickering in the texture of clothing wrinkles, and MuseV [39] struggles with generating consistent text on clothing. In contrast, our method maintains stable inter-frame differences without obvious artifacts, demonstrating better temporal smoothness in our generation results. Videos are included on the project page. This enhancement in temporal smoothness is likely due to the robustness provided by our confidence-aware pose guidance, which effectively mitigates the impact of inaccurate pose inputs and temporal noise. By intelligently weighting the influence of pose signals based on their confidence, our method ensures that the generated videos maintain a high level of temporal smoothness in the presence of noise.
@@
In Table 8, we present a quantitative comparison of our method against state-of-the-art approaches using the FID-VID [41] and FVD [42] metrics on the test sequences from the TikTok [40] dataset. The results indicate that our method achieves a notable performance advantage over all existing methods in terms of both metrics.
@@
To supplement our quantitative and qualitative evaluations, we conduct a user study to assess the subjective preferences of participants regarding the generated videos on the TikTok dataset test split. The study involves showing two video clips—one generated by our method and the other by one of the baseline methods—to a diverse group of users. Participants are instructed to select the video that they perceived as having higher quality, considering factors such as image quality, flickering, and the temporal smoothness of characters and clothing. We collected data from 36 participants, with each participant evaluating 6 video pairs for our method against each baseline method. As shown in Figure 8, the results indicate a strong preference for MimicMotion over the baseline methods. In comparison to MagicPose and Moore, the participants almost favored all videos produced by our method. Despite MuseV showing higher image quality compared to other baselines, the preference for videos produced by our method still reached 75.5%. These findings align with our qualitative and quantitative evaluation, reinforcing the effectiveness of our method in meeting user expectations for high-quality human video generation.
@@
[…]
MagicPose [4] 13.3 916
Moore [38] 12.4 728
MuseV [39] 14.6 754
MimicMotion (ours) 9.3 594
Figure 7: Quantitative comparison to state-of-the-art methods. We evaluate FID-VID and FVD on the TikTok dataset test split. MimicMotion is the best in both metrics.
[…]
Figure 8: User study. Preference of MimicMotion (ours) over baseline methods on the TikTok dataset test split. Users prefer MimicMotion over other methods.
4.3 Ablation Study
@@
Figure 9 shows the effectiveness of confidence-aware pose guidance. Each row corresponds to one example. On the left side, we show three images used to extract the pose. On the right side, we plot the guiding signals corresponding to the pose estimation, both with and without confidence-aware pose guiding. From the guiding signals, we can see that there are errors in the pose estimated by DWPose. Nevertheless, our confidence-aware design minimizes the impact of incorrect pose estimation in guidance signals.
@@
In contrast, by integrating confidence scores into the pose representation, our method effectively mitigates these issues. The confidence scores provide a measure of reliability for each keypoint, allowing the system to weigh the guidance signals accordingly. Specifically, keypoint with lower confidence, which typically correspond to inaccurate keypoints caused by occlusion or motion blur, will be of less significance in the guidance. This approach leads to clearer and richer pose guidance, as the influence of potentially erroneous keypoints is reduced. The corresponding generation results demonstrate how our method enhances the robustness of generation against false guiding signals (Pose 1 and Pose 2) and offers visibility hints to resolve the front-back ambiguity of 2D pose estimation (Pose 3).
Figure 9: Confidence-aware pose guiding. This design enhances generation robustness to false guiding signals (Pose 1&2) and provides visibility hints to tackle pose ambiguity (Pose 3).
@@
In conjunction with confidence-aware pose guiding, we further improve the quality of hand generation by assigning a higher weight to the hand region in the training loss. Figure 10 compares the generation result with and without hand region enhancement, using the same reference image and pose guidance. All experiments incorporate confidence-aware pose guidance. The hands in the first row are cropped from the generated video frames of a model trained without hand region enhancement, which exhibits noticeable distortions, such as irregular and misplaced fingers. In contrast, the results of the model trained with hand region enhancement (the second row) show consistent improvements in hand generation quality and a reduction in hand distortion. These results show the effectiveness of the proposed hand region enhancement design, which substantially mitigates hand distortion, which is a prevalent challenge in diffusion-based models.
Moreover, hand region enhancement improves the visual appeal of the generated content. The hand region is often the area of interest that human observers tend to focus on. By emphasizing the hand regions, we align the regional preferences of the training process with human preferences, thereby enhancing the visual appeal of the generation results.
Figure 10: Hand region enhancement. Training with hand enhancement consistently reduces hand distortion and improves visual appeal, using the same reference image and pose guidance.
@@
To achieve seamless transitions between video segments, we introduce progressive latent fusion, a technique that gradually blends frames in the overlapped regions of consecutive video segments. The original MultiDiffusion approach employs a simple averaging of frames within the overlap region. As illustrated in Figure 11(a), this method assigns equal weight to all frames in the overlap region, irrespective of their temporal position (whether they are closer to the preceding or subsequent segment). This lack of a gradual transition in influence from one segment to another can cause abrupt transitions and noticeable flickers in the video. This is evident in the y-t slice shown on the left, where the transition at segment boundaries is abrupt. The right side of the figure shows four frames at the segment boundary. Note that the background in the top-left corner (enlarged) is initially clear in segment 1. It suddenly becomes blurry in the overlapped region and then suddenly reverts to clearer in the main part of segment 2. This artifact is not observed when progressive latent fusion is applied.
The proposed progressive latent fusion approach (see Figure 11(b)) effectively mitigates these issues. The y-t slice on the left demonstrates that this method enables a smooth transition across segment boundaries, eliminating the abrupt changes seen in the original approach. The right side of the figure demonstrates the relevance of sudden blurring. This strategy significantly mitigates flicking artifacts, thus improving the overall visual temporal coherence for long video generation.
@@
Figure 11: Progressive latent fusion enables smooth segment (clip) transitions for long video generation. (a) Simple averaging in overlapped regions causes temporal discontinuity, appearing as vertical strips in the Y-T slice. (b) Progressive latent fusion enables smooth transitions.

### v2

@@ 3.4 Confidence-Aware Pose Guidance
Inaccurate pose estimation has a negative impact on the model’s training and inference. However, accurate pose estimation is challenging in dynamic videos. The limited capability of the pose estimation model, like DWPose (Yang et al., 2023), is only one aspect of the reason. The more significant cause is the inherent uncertainty of pose from dynamic appearances and motions. Specifically, incorrect pose guidance signals can mislead the model, resulting in the generation of inaccurate or distorted outputs, as illustrated in Figure 7. Moreover, noisy pose guidance signals can lead to overfitting on samples with incorrect poses, potentially causing training instability. This in turn may hinder the model’s training.
For this problem, we propose confidence-aware pose guidance, which leverages the confidence scores associated with each keypoint from the pose estimation model. These scores reflect the likelihood of accurate detection, with higher values indicating higher visibility, less self-occlusion from other body parts, and motion blur. Instead of applying a fixed confidence threshold to filter the keypoints, as commonly adopted in prior works (moo, 2024; Chang et al., 2023), we utilize brightness of the keypoints and limb to represent the confidence level of pose estimation. Specifically, we multiply the color assigned to each keypoint and limb by its confidence score. Consequently, keypoints and corresponding limbs with higher confidence scores will appear more significant on the pose guidance map. This method enables the model to prioritize more reliable pose information in its guidance, thereby enhancing the overall accuracy of pose-guided generation. In this way, the uncertainty of pose estimation can be conveyed through the pose guidance, making pose guidance more informative.
@@
Moreover, we employ pose estimation and the associated confidence scores to alleviate region-specific artifacts, such as hand distortion, which are prevalent in the diffusion-based image and video generation models. Specifically, we identify reliable regions via thresholding keypoint confidence scores. By setting a threshold, we can distinguish between keypoints that are confidently detected and those that may be unreliable due to self-occlusion or motion blur. Keypoints with confidence scores above the threshold are considered reliable. We implement a masking strategy that generates masks based on a confidence threshold. We unmask areas where confidence scores surpass a predefined threshold, thereby identifying reliable regions. When computing the loss of the video diffusion model, the loss values corresponding to the unmasked regions are amplified by a certain scale so they can have more effect on the model training than other masked regions.
@@
Specifically, to mitigate hand distortion, we compute masks using a confidence threshold for keypoints in the hand region. Only hands with all keypoint confidence scores exceeding this threshold are considered reliable, as a higher score correlates to higher visual quality. We then construct a bounding box around the hand by padding the boundary of these keypoints, and the enclosed rectangle is designated as unmasked. This region is subsequently assigned a larger weight in the loss calculation during the training of the video diffusion model. This selective unmasking and weighting process biases the model’s learning towards hands, especially hands with higher visual quality, effectively reducing distortion and improving the overall quality.
3.5 Progressive Latent Fusion for Long Videos
@@
Figure 3: Overview of our approach for long video generation. The colored boxes represent latent video frames. A darker color means a higher weight. The dashed boxes represent video frame features involved in latent fusion.
@@
Limited by computation resources, generating long videos containing a large number of frames is challenging. For this problem, prior works like MultiDiffusion (Bar-Tal et al., 2023) successfully use latent fusion for panoramic image generation. A similar idea can be applied to the video generation task. A straightforward approach is directly applying MultiDiffusion in the time domain, as in Lumiere (Bar-Tal et al., 2024). However, unlike spatial discontinuities between image tiles, temporal discontinuities are more noticeable to viewers, resulting in flickering or abrupt content changes. We address this by a progressive approach for generating long videos with high temporal continuity.
@@
Progressive latent fusion is training-free and is integrated into the denoising process of the latent diffusion model during inference. Figure 3 shows an overview of this process. We omit the VAE for brevity. The denoising process is done in latent space in our method. In general, there are $T$ denoising steps in total and our latent fusion is applied within each step. For a long given pose sequence, we use a pre-defined strategy for splitting the whole sequence into segments, consisting of a fixed number of frames per segment (denoted as $N$), with a certain number ($C$) of overlapped frames between every two adjacent segments. For the sake of generation efficiency, it is common to assume that $C\ll N$. During each denoising step, video segments are firstly denoised separately with the trained model, conditioning on the same reference image and the corresponding sub-sequence of poses. Algorithm 1 shows the specific details of progressive latent fusion. As inputs, the reference image is denoted as $I_{\text{ref}}$, and the pose frame corresponding to $j$-th frame in $i$-th video segment is denoted as $P_{i}^{j}$. We use $\mathbf{z}_{i}^{j}$ to denote the latent feature of $j$-th frame in $i$-th video segment. The denoising process starts from the maximum time step $T$ and the latent features are initialized with a normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$. Within each denoising step at time step $t$, the reversed diffusion process defined by the trained model (DM) is applied to the latent features of each video segment numbered $i$ separately, with $\mathbf{z}_{i}$, $I_{\text{ref}}$, $P_{i}$ and $t$ as inputs. During the latent fusion stage, for every two adjacent video segments, the involved video frames are then fused. To avoid the corruption of temporal smoothness near video segment boundaries after latent fusion, we propose progressive latent fusion. For a video frame involved in latent fusion, its fusion weight is determined by its relative position in the video segment it belongs. Specifically, if a frame is close to the segment it belongs to, it will be assigned a heavier weight. For implementation, a fusion scale is pre-defined as $\lambda_{\text{fusion}}=1/(C+1)$ for controlling the level of latent fusion.
Algorithm 1 Progressive latent fusion for long videos.
1:  Input:
2:  $I_{\text{ref}}$: Reference image;
3:  $P_{i}^{j}$: Pose of $j$-th frame in $i$-th video segment;
4:  $\mathbf{z}_{i}^{j}$: The latent of $j$-th frame in $i$-th video segment;
5:  $N$: The number of frames in each video segment;
6:  $C$: The number of overlapped frames.
7:  Output: A long sequence of latent video frames.
8:  $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
9:  $\lambda_{\text{fusion}}\leftarrow 1/(C+1)$
10:  for $t=T$ to $1$ do
11:     for $i=1,2,\dots$ do
12:        $\mathbf{z}_{i}\leftarrow\text{DM}(\mathbf{z}_{i},I_{\text{ref}},P_{i},t)$
13:     end for
14:     for $i=1,2,\dots$ do
15:        for $j=1$ to $N$ do
16:           if $i>1$ and $j\leq C$ then
17:              $\mathbf{z}_{i}^{j}\leftarrow j\lambda_{\text{fusion}}\mathbf{z}_{i}^{j}+(1-j\lambda_{\text{fusion}})\mathbf{z}_{i-1}^{N-C+j}$
18:           else if $j>N-C$ then
19:              $\mathbf{z}_{i}^{j}\leftarrow(N+1-j)\lambda_{\text{fusion}}\mathbf{z}_{i}^{j}+(C-N+j)\lambda_{\text{fusion}}\mathbf{z}_{i+1}^{C-N+j}$
20:           end if
21:        end for
22:     end for
23:  end for
24:  return $\{\mathbf{z}_{i}^{1:N-C}\}_{i=1,2,\cdots}$
After applying $T$ iterative denoising steps, we aggregate the final long sequence by combining the denoised overlapped video segments in latent space. The resulting aggregated long video in latent space is then fed to the temporal VAE decoder to reconstruct the long video.
@@ 4.1 Implementation Details
Due to the lack of suitable high-quality academic datasets, we collect 4,436 human dancing videos from the internet for training. The average length is 20.1s. We adopt the pre-trained weights from SVD (Blattmann et al., 2023). The PoseNet is trained from scratch. We train our model on 8 NVIDIA A100 GPUs for 20 epochs, with a batch size of 8 and 16 frames per clip. The loss weight of the hand region is 10. The learning rate is $10^{-5}$ with a linear warmup of 500 iterations. We tune all parameters in the UNet and PoseNet.
@@ 4.2 Comparison to State-of-the-Art Methods
We compare our method with latest state-of-the-art pose-guided human video generation methods, including MagicAnymate (Xu et al., 2024), MagicPose (Chang et al., 2023), Moore-AnymateAnyone (moo, 2024), and MuseV (Xia et al., 2024). For the testing protocol of previous works (Wang et al., 2023; Chang et al., 2023), we adopt the TikTok (Jafarian & Park, 2021) dataset and use sequence 335 to 340 for our evaluation.
We conduct qualitative and quantitative comparisons, complemented by a user study. Each method has a different input aspect ratio. To ensure a fair comparison, we only consider the central square region of the videos. Specifically, we individually apply a center crop to the reference image and pose sequence to match each method’s input aspect ratio. Then, we extract the center squares from the generated videos for comparison across different methods.
@@
Figure 4: Qualitative comparison to the state-of-the-art methods on TikTok test set. Our method achieves better hand generation quality and adheres better to the reference pose. Note that our method is not trained on the TikTok dataset.
@@
Importantly, our method shows superior temporal smoothness, characterized by smooth motion and minimal flickering. Figure 5 demonstrates this through pixel-wise differences between consecutive frames. While MagicPose (Chang et al., 2023) shows abrupt transitions, Moore-AnymateAnyone (moo, 2024) exhibits unstable clothing textures, and MuseV (Xia et al., 2024) struggles with consistent text rendering, our method maintains stable frame-to-frame transitions without visible artifacts. Videos in the supplementary material further demonstrate this improvement. This enhancement is enabled by our confidence-aware pose guidance, which mitigates the impact of inaccurate pose inputs and the associated temporal noise. By weighting pose signals by confidence, our method ensures high temporal smoothness in the presence of pose guidance noise.
Figure 5: Comparison of temporal smoothness via frame differences. MagicPose exhibits abrupt transitions, whereas Moore and MuseV show flickering texture and text. In contrast, our method maintains stable frame differences.
@@
Our method outperforms existing approaches across all evaluation metrics, as shown in Table 1. We evaluated performance on test sequences from the TikTok dataset (Jafarian & Park, 2021) using both video-level metrics FID-VID (Balaji et al., 2019) and FVD (Unterthiner et al., 2018), as well as frame-level metrics PSNR and SSIM (Wang et al., 2004) following (Wang et al., 2023). The results demonstrate our method’s superior performance compared to current state-of-the-art approaches.
[…]
MagicAnymate 16.2 848 0.740 17.5
MagicPose 13.3 916 0.776 18.8
Moore 12.4 728 0.758 18.7
MuseV 14.6 754 0.766 17.6
MimicMotion (ours) 9.3 594 0.795 20.1
Table 1: Quantitative comparison to state-of-the-art methods: MagicAnymate (Xu et al., 2024), MagicPose (Chang et al., 2023), Moore (moo, 2024), MuseV (Xia et al., 2024). We evaluate on the TikTok dataset test split. MimicMotion is the best in all metrics.
@@
To supplement our quantitative and qualitative evaluations, we conduct a user study to assess the subjective preferences of participants regarding the generated videos on the TikTok dataset test split. The study involves showing two video clips—one generated by our method and the other by one of the baseline methods—to a diverse group of users. Participants are instructed to select the video that they perceived as having higher quality, considering factors including image quality and temporal smoothness of characters and clothing. We collected data from 36 participants, with each participant evaluating 6 video pairs for our method against each baseline method. As shown in Figure 6, the results indicate a strong preference for MimicMotion over the baseline methods. In comparison to MagicPose and Moore, the participants almost favored all videos produced by our method. Despite MuseV showing higher image quality compared to other baselines, the preference for videos produced by our method still reached 75.5%. These findings align with our qualitative and quantitative evaluation, reinforcing the effectiveness of our method in meeting user expectations for high-quality human video generation.
Figure 6: Preference of videos generated by MimicMotion (ours) over baseline methods on the TikTok dataset test split. Users consistently prefer MimicMotion over other methods.
4.3 Ablation Study
As shown in Table 2, confidence-aware pose guiding and hand region enhancement improve all metrics. Progressive latent fusion notably improves FVD scores, indicating better temporal coherence, though other metrics may not reflect this improvement. We analyze each component below.
@@ 4.3 Ablation Study
14.6 776 0.760 18.0
$\checkmark$ 15.0 678 0.758 17.9
$\checkmark$ $\checkmark$ 12.2 623 0.787 18.4
$\checkmark$ $\checkmark$ $\checkmark$ 9.3 594 0.795 20.1
Table 2: Ablation studies of hand region augmentation (hand), confidence-aware pose guiding (conf.), and progressive latent fusion (prog.).
@@
Figure 7 shows the effectiveness of confidence-aware pose guidance. Each row corresponds to one sample. The left side shows three images to extract the pose. On the right side, we plot the corresponding pose guiding signals, both with and without confidence-aware pose guiding. From the guiding signals, we can see that there are errors in the pose estimated by DWPose. Nevertheless, our confidence-aware design minimizes the impact of incorrect pose estimation in guidance signals.
@@
In contrast, by integrating confidence scores into pose guidance, our method effectively mitigates these issues. The confidence scores provide a measure of reliability for each keypoint, allowing our model to weigh the guidance signals accordingly. Specifically, keypoint with lower confidence, which typically correspond to inaccurate keypoints caused by self-occlusion or motion blur, will be of less significance in the guidance. This approach leads to clearer and richer pose guidance, as the influence of potentially erroneous keypoints is reduced. The corresponding generation results demonstrate how our method enhances the robustness of generation against false guiding signals (Pose 1 and Pose 2) and offers visibility hints to resolve the front-back ambiguity of 2D pose estimation (Pose 3).
Figure 7: Confidence-aware pose guiding enhances generation robustness to false guiding signals (Pose 1&2) and provides hints to handle self-occlusion (Pose 3).
@@
In conjunction with confidence-aware pose guiding, we further improve hand generation quality by assigning a higher weight to the hand region in the training loss. Figure 8 compares the generation result with and without hand region enhancement, using the same reference image and pose guidance. All experiments incorporate confidence-aware pose guidance. The hands in the first row are cropped from the generated video frames of a model trained without hand region enhancement, which exhibits noticeable distortions, such as irregular and misplaced fingers. In contrast, the results of the model trained with hand region enhancement (the second row) show consistent improvements in hand generation quality and a reduction in hand distortion. These results show the effectiveness of the proposed hand region enhancement design, which substantially mitigates hand distortion, which is a prevalent challenge in diffusion-based models. As humans often focus on hands, this design aligns the training process with human preferences, thereby enhancing the overall visual quality of the results. There is no noticeable degradation in the quality of the remaining parts.
Figure 8: Hand region enhancement consistently reduces hand distortion and improves visual appeal, using the same reference image and pose guidance.
@@
To achieve seamless transitions between video segments, we introduce progressive latent fusion, a technique that gradually blends frames in the overlapped regions of consecutive video segments. Figure 9 presents a comparison of a sample transition with an overlap of 6 frames between two 16-frame segments. We visualize 4 sample frames. The original MultiDiffusion approach employs a simple averaging of frames within the overlap region. This method assigns equal weight to all frames in the overlap region, irrespective of their temporal position (whether they are closer to the preceding or subsequent segment). It lacks a gradual transition from one segment to another, causing noticeable flickers at denoising window boundaries. This is evident in the y-t slice shown on the left of Figure 9(a), where the transition at segment boundaries is abrupt. The right side of the figure shows four frames at the segment boundary. Note that the background in the top-left corner (enlarged) is initially clear in segment 1. It suddenly becomes blurry in the overlapped region and then suddenly reverts to clearer in the main part of segment 2. This artifact is not observed when progressive latent fusion is applied.
The proposed progressive latent fusion approach (see Figure 9(b)) mitigates these issues. The left y-t slice shows smooth transitions across segment boundaries, avoiding abrupt changes. The right side of the figure demonstrates that this approach eliminates a sudden blurring, mitigating flicking artifacts, thus improving the overall visual temporal coherence for long video generation.
@@
Figure 9: Progressive latent fusion ensures smooth segment transitions by avoiding temporal discontinuities of weights in simple averaging, which reduces artifacts in Y-T slices.